OpenStudy (anonymous):

4x*2-6y=14 slope intersept form

OpenStudy (anonymous):

help dina

OpenStudy (lynfran):

4x^2-14=6y 6y=4x^2-14 y=(4x^2-14)/6

OpenStudy (lynfran):

can also be written y=4x^2/6 - 14/6 which can also be simplify

OpenStudy (lynfran):

Do u understand?

OpenStudy (imtiaz7):

As @LynFran say Which can also be simplify as: y=4x^2/6 - 14/6 y=2x^2/3 - 7/3 ( It comes in slope intercept form that is y=mx+b)